AGW2 Alignment with LUCI2@RFBG 20151025 UT

Observers: JHill, DMiller, GTaylor, DThompson (Tucson)
Telescope Operator: GBechetti (LBTO)
AO Support: DMiller
Telescope Support: MWagner (LBTO)
Software Support: TSargent (Tucson)
Instrument Support: ESolheid (LBTO)
Engineering Support: DAshby

Summary

The first 2 hours were lost to humidity and wind. When we did open, the seeing was really awful (we saw an 11.8 on the DIMM, not believable but impressive). We verified that the guiding hotspot is vaguely correct, but it isn't possible to make a precise tuning in this seeing. It is also not possible to check the field aberrations.

What we did accomplish was adjusting the DX M3 tip-tilt to minimize the pupil wobble (1/4 subap), and we adjusted the nominal pupil position (WFScam_hotspot).

We turned the telescope over to Dave T for LUCI2 closed-dome calibration at 5:45 UT.

Preparations

The rigid (only) secondary was moved from the SX to DX side this afternoon.

Details

The evening starts with humidity and wind both near threshold. Clouds are approaching from BOTH the southeast hurricane patricia and from the northwest.

Open

03:00 - Open in a fairly clear sky. T=3.0 D=-1.9 Wind 15m/s from the East.

Initial DX M2 Alignment

03:10 Seeing is really bad = 4-5 arcsec. Defocussed pupil does not have obvious coma. Pointing was off by ~30 arcsec. We didn't make any estimated binodal adjustment from the normal AdSec collimation model. Focus Global Offset eyeballed at -3.0 mm

AGW2 Guide Hotspot Alignment

03:17 ACTIVE preset to BS9188 - block Zernike reporting, but we didn't see anything.

03:20 ACTIVE preset to ACT0338 - trying to divine the focus by eye on overfilled wfs. Adopted focus global offset -3.4 mm

03:43 ACTIVE preset to AO54 (8.9 mag) - Our focus is correct within +-0.05 mm, but active optics is hopeless in seeing this bad.

03:53 unblock the Zernikes (rms wavefront error is 666 nm!) - This removed some coma (M2 Active Y=+1.5 mm), but didn't change the focus. Seeing continues to bounce between 3-5 arcsec as measured by guider FWHM. DIMM is recording readings between 3.5 and 11.8 arcsec. We can say that the hotspot is vaguely in the correct position.

AGW2 Pupil Wobble

Starting at DX M3 Absolute Tip=-1200 Tilt=250. The peculiar thing is that the PSF collimation file has Global Offsets of Tip=-200 Tilt=+250. Take a wfs image at PA=0 and PA=180 deg.

04:18 Reduce Tilt by 250 arcsec to 0 absolute.

04:22 Increase Tip by 200 arcsec to -1000 absolute.

04:27 Increase Tip by 200 arcsec to -800 absolute. This is improved.

04:40 Decrease Tip by 100 arcsec to -900 absolute.

04:42 Increase Tip by 100 arcsec to +100 absolute.

Adopted DX M3 position is absolute Tip=-900 Tilt=+100 at nominal -25.725 deg selector. A set of 4 PA images concludes that the radius of the residual pupil wobble is about 1/4 subap. This is a successful adjustment, although it wasn't far off to start with.

For this new M3 Tip-Tilt the pointing is IE=-7 CA=-179 . I did NOT update the pointing model zeropoints.

* New values for WFScam_hotspot_x,y are 378,265 pixels* in LUCI_R.cfg.

Updated in OSS/positions.conf
oss.side[1].terc.focal_stations.bentGregorianFront              map  0:-900.0   1:100.00 2:0.0  3:-25.727
oss.side[1].terc.focal_stations.bentGregorianArgos              map  0:-900.00  1:100.00 2:0.0  3:-25.734

Changed global offsets to 0,0 in DXTMLUCIFERCollimation.dat.

05:08 ACTIVE preset to another field in the South to see if the seeing is any better over there......

Stop OSS and GCS to ingest the changes. Initialize the PSF M3 lookup table.

05:30 resend the preset.

05:50 Handover to DaveT for LUCI2 closed-dome calibrations.

AGW2 Aligning Probe to Rotator Center

-- %USERSIG{JohnHill - 2015-10-25}%
Topic revision: r3 - 25 Oct 2015, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback